An Aluminium bar is found to measure 50cm at 20°C when measured with brass scale correct at 0°C . what is the true length of the bar at  20 °C?
          

Grade:9

1 Answers

Shubhangi Shilendra
5 years ago
When temperature increases, length of obj observed by scale is less than correct length.
hence true value at measured temperature 20 degrees Celsius is
Lt= true value of length. Lm= length measured at 20 degrees celsius
a= linear coefficient of exp of brass= 18x10^-6
Lt= Lm+ (Lm x a x change in temperature)
   =20 + (20 x 18 x 10^-6 x 20)
   = 20 + (72 x 10^-4)
   =20.0072cm